# SecureFalcon: Are We There Yet in Automated Software Vulnerability Detection with LLMs?

### v2

@@ VI-C Comparison of SecureFalcon
Table XII compares various machine learning (ML) models and their accuracy in performing multiclass and binary classification tasks on the FormAI and FalconVulnDB datasets. Large Language Models (LLMs) such as RoBERTa, BERT, CodeBERT, and SecureFalcon show a notable performance difference compared to traditional ML models for multiclass classification. SecureFalcon, in particular, stands out with the highest accuracy of 0.92 on both datasets, significantly outperforming other LLM models like RoBERTa and BERT, which achieve accuracies of around 0.70-0.85. Among the traditional ML models, Random Forest (RF) performs the best with accuracies of 0.77 on the FormAI dataset and 0.81 on the FalconVulnDB dataset, although it still lags behind the LLM models.
@@
TABLE XII: Comparison of SecureFalcon with LLM models and traditional machine learning using the FormAI dataset and FalconVulnDB dataset.
Task ML type Model Accuracy
FormAI dataset FalconVulnDB dataset
Multiclass Classification LLM models RoBERTa 0.70 0.85
BERT 0.70 0.85
CodeBERT 0.71 0.88
SecureFalcon 0.92 0.92
Traditional ML models KNN 0.76 0.80
LR 0.73 0.77
NB 0.64 0.68
SVM 0.73 0.77
RF 0.77 0.81
DT 0.72 0.60
LDA 0.73 0.77
Binary Classification LLM models RoBERTa 0.82 0.86
BERT 0.81 0.87
CodeBERT 0.83 0.89
SecureFalcon 0.94 0.92
Traditional ML models KNN 0.76 0.86
LR 0.86 0.86
NB 0.76 0.81
SVM 0.86 0.89
RF 0.91 0.89
DT 0.87 0.86
LDA 0.86 0.85
@@
VI-D Future Works
@@
As future work, SecureFalcon, a fine-tuned Large Language Model (LLM) for software vulnerability detection, aims to enhance its capabilities by providing comprehensive explanations to developers. These enhancements include generating detailed reports describing detected vulnerabilities, highlighting the affected code snippets, and analyzing the potential impact. Furthermore, SecureFalcon will outline steps to reproduce the issues, including environment setup, input data, and execution steps, and offer remediation guidance with fix recommendations and secure coding practices. By integrating these features with development tools, SecureFalcon will ensure real-time feedback and automated code reviews, significantly aiding developers in understanding, reproducing, and resolving security issues effectively.

### v3

@@ VI-D Comparison of SecureFalcon with LLM, ML, and DL models
Table XIII compares various machine learning (ML) models and their accuracy in performing multiclass and binary classification tasks on the FormAI and FalconVulnDB datasets. Large Language Models (LLMs) such as RoBERTa, BERT, CodeBERT, and SecureFalcon show a notable performance difference compared to traditional ML models for multiclass classification. SecureFalcon, in particular, stands out with the highest accuracy of 0.92 on both datasets, significantly outperforming other LLM models like RoBERTa and BERT, which achieve accuracies of around 0.70-0.85. Among the traditional ML models, Random Forest (RF) performs the best with accuracies of 0.77 on the FormAI dataset and 0.81 on the FalconVulnDB dataset. However, it still lags behind the LLM models.
@@
Can traditional machine learning models effectively capture vulnerability patterns without sophisticated semantic understanding? For the FormAI dataset, which is synthetic and constructed in a manner favoring simpler patterns, the answer appears to be yes. In the multiclass scenario on the FormAI dataset, CodeBERT—a general pre-trained LLM fine-tuned for classification—achieves an accuracy of 0.71, while specific traditional ML models, such as Random Forest (RF), reach up to 0.77. This suggests that vulnerability patterns in FormAI can be successfully modeled using classical feature engineering or straightforward pattern recognition. Traditional ML methods, often relying on syntactic features (e.g., token frequencies or specific keywords), perform competitively because well-established, surface-level indicators may represent the dataset’s vulnerabilities.
But what happens when the complexity of vulnerabilities increases, demanding deeper semantic comprehension? In the FalconVulnDB dataset, where vulnerabilities involve more subtle and context-dependent cues, traditional ML models show a performance plateau (e.g., RF at 0.81 accuracy in multiclass classification). In contrast, LLM-based models—pre-trained on extensive corpora and adept at contextualizing code semantics—excel (e.g., CodeBERT at 0.88). The heightened complexity and nuance of FalconVulnDB vulnerabilities seemingly require the richer representational capacity and more advanced reasoning capabilities that LLMs provide.
Table XVI compares our proposed SecureFalcon model against several state-of-the-art deep learning-based vulnerability detection methods, including SySeVR [41], VulDeePecker [6], and Devign [42]. While the earlier approaches often rely on hand-crafted features (e.g., syntax characteristics or code gadgets) and are typically limited to detecting a narrower set of vulnerabilities, SecureFalcon leverages a large language model architecture derived from Falcon-40B. This allows SecureFalcon to incorporate both synthetically generated data and real-world vulnerability instances drawn from a broad set of sources (FalconVulnDB), ultimately enabling it to target a far more comprehensive range of CWE types, including the industry-recognized Top 25 Most Dangerous CWEs  [^11] https://cwe.mitre.org/top25/archive/2024/2024_cwe_top25.html. In contrast, the earlier methods examined focus on a smaller subset of vulnerabilities and do not integrate synthetic data to enhance coverage and robustness.
TABLE XIII: Comparison of SecureFalcon with LLM models and traditional machine learning using the FormAI and FalconVulnDB datasets.
Task ML type Model FormAI dataset FalconVulnDB dataset
Acc Prec Rec F1 Acc Prec Rec F1
Multiclass LLM models RoBERTa 0.70 0.72 0.68 0.70 0.85 0.87 0.84 0.85
BERT 0.70 0.71 0.69 0.70 0.85 0.86 0.84 0.85
CodeBERT 0.71 0.73 0.70 0.71 0.88 0.89 0.87 0.88
SecureFalcon 0.92 0.92 0.93 0.93 0.92 0.91 0.91 0.90
Traditional ML KNN 0.76 0.75 0.76 0.76 0.80 0.79 0.80 0.80
LR 0.73 0.74 0.72 0.73 0.77 0.78 0.76 0.77
NB 0.64 0.66 0.62 0.64 0.68 0.69 0.67 0.68
SVM 0.73 0.73 0.72 0.73 0.77 0.78 0.77 0.77
RF 0.77 0.78 0.76 0.77 0.81 0.81 0.80 0.81
DT 0.72 0.72 0.71 0.72 0.60 0.61 0.59 0.60
LDA 0.73 0.74 0.73 0.74 0.77 0.78 0.76 0.77
Binary LLM models RoBERTa 0.82 0.83 0.81 0.82 0.86 0.87 0.85 0.86
BERT 0.81 0.82 0.80 0.81 0.87 0.88 0.86 0.87
CodeBERT 0.83 0.84 0.82 0.83 0.89 0.90 0.88 0.89
SecureFalcon 0.94 0.94 0.94 0.94 0.92 0.93 0.91 0.92
Traditional ML KNN 0.76 0.77 0.75 0.76 0.86 0.86 0.85 0.86
LR 0.86 0.87 0.86 0.86 0.86 0.87 0.85 0.86
NB 0.76 0.77 0.74 0.75 0.81 0.82 0.80 0.81
SVM 0.86 0.87 0.85 0.86 0.89 0.90 0.88 0.89
RF 0.91 0.91 0.90 0.91 0.89 0.90 0.88 0.89
DT 0.87 0.88 0.85 0.86 0.86 0.87 0.84 0.85
LDA 0.86 0.86 0.85 0.85 0.85 0.85 0.85 0.85
@@
TABLE XIV: Comparison of SecureFalcon with C/C++ Vulnerability Detection Tools
Tool/Model Accuracy (%) Precision (%) Recall (%) F1 Score (%) Inference Time (ms/sample)
SecureFalcon 93.5 92.0 91.8 91.9 500
Cppcheck 81.0 79.5 80.2 79.8 150
Clang Static Analyzer 85.2 83.0 84.0 83.5 200
ESBMC 90.0 91.0 87.5 89.2 2000
VI-E Comparison of SecureFalcon with C/C++ Vulnerability Detection Tools
Table XIV presents a comparative evaluation of SecureFalcon, Cppcheck  [^12] https://cppcheck.sourceforge.io/, Clang Static Analyzer [^13] https://clang-analyzer.llvm.org/, and ESBMC  [^14] https://github.com/esbmc/esbmc. We randomly sampled 50 code segments from the test part of our FalconVulnDB dataset and manually validated each prediction. For the inference of SecureFalcon, we used an NVIDIA A100 GPU with 40 GB of memory. While static analyzers like Cppcheck and Clang Static Analyzer excel in speed and are easy to integrate, they often miss context-dependent vulnerabilities. ESBMC is a mature bounded model checker that supports the verification of single- and multi-threaded C/C++ programs, but it has limitations: it is both slower and requires compilable code. Some code snippets in the dataset—particularly those that are not fully compilable—could not be analyzed by ESBMC, thereby reducing its practical coverage. In contrast, SecureFalcon’s LLM-based approach not only offers strong detection rates and nuanced semantic understanding but can also handle snippets that do not compile, expanding its applicability. Although SecureFalcon’s inference time is longer than lightweight tools, the accuracy gains and flexibility in handling diverse code samples justify the trade-off in scenarios demanding thorough security checks.
Ultimately, while traditional static analyzers are suitable for quick preliminary scans, ESBMC is valuable for in-depth formal verification of compilable code. SecureFalcon is a versatile solution that effectively balances accuracy, context awareness, scalability, and broader code coverage. Furthermore, combining traditional static analyzers and ESBMC with LLMs like SecureFalcon can provide synergistic benefits, leveraging the speed and simplicity of static analyzers, the rigorous verification of formal tools, and the deep contextual understanding of LLMs. This integrated approach ensures comprehensive vulnerability detection, maximizing both efficiency and accuracy across various development and deployment scenarios.
TABLE XV: Ablation results showing the impact of removing or modifying key components and configurations of SecureFalcon. Each setting was evaluated using the binary classification task on the FormAI dataset (LR = 2e-5). The baseline configuration refers to all features and parameters as listed in Table IV.
Configuration Description of Change Accuracy F1-score
Baseline (Full Model) All components enabled: positional encodings (514 max), special token embeddings, 12 attention heads, dropout=0.1 0.94 0.94
No Positional Embeddings Removed positional embeddings from the transformer layers, disabling positional indexing of tokens 0.91 (-0.03) 0.90 (-0.04)
No Special Token Embeddings Excluded the domain-specific tokens (keywords, punctuation, API calls) from the vocabulary 0.89 (-0.05) 0.91 (-0.03)
Fewer Attention Heads Reduced number of attention heads from 12 to 6, decreasing model’s representational capacity 0.90 (-0.04) 0.90 (-0.04)
Increased Dropout Doubled attention and hidden dropout from 0.1 to 0.2, increasing regularization 0.92 (-0.02) 0.93 (-0.01)
TABLE XVI: Comparison of SecureFalcon with deep learning-based vulnerability detection methods.
Model Method Datasets Used Synthetic Data Real Data CWE Types Targeted Top 25 CWEs
SySeVR [41] Syntax & semantics-based vector representations 4 kinds of vulnerability syntax characteristics No Yes Limited (fewer vulnerabilities) No
VulDeePecker [6] Code gadgets & deep learning classification Curated dataset from known software No Yes Limited (e.g., buffer overflow) No
Devign [42] Graph neural networks on code semantics 4 large-scale open-source C projects No Yes Limited (depends on projects) No
SecureFalcon LLM-based (derived from Falcon-40B) FalconVulnDB (combining multiple datasets) Yes Yes Broad (Top 25 + beyond) Yes
VI-F Inference Time of SecureFalcon
Inference performance is a critical metric for evaluating the applicability of machine learning models in real-world scenarios, particularly for tasks such as vulnerability classification in C/C++ codebases.SecureFalcon, with its compact architecture of 121 million parameters, demonstrates exceptional efficiency with a Time to First Token (TTFT) of 0.3 seconds and a total classification time of 0.6 seconds for typical input sizes of 300-500 tokens. It achieves a high throughput of 45-50 requests per second (RPS) and processes 95 tokens per second on a single NVIDIA A100 GPU. Furthermore,SecureFalcon is optimized for CPU inference, achieving competitive latency and throughput on hardware such as the AMD Ryzen 7 7840U CPU (8 cores, 16 threads, 30W TDP) with 32GB LPDDR5X RAM. These results highlight SecureFalcon’s suitability for edge devices and on-premises deployments, offering the speed and scalability required for enterprise-scale vulnerability detection. By delivering low-latency and high-throughput inference across a range of hardware configurations,SecureFalcon enables seamless integration into real-time workflows, including CI/CD pipelines, and supports efficient analysis of large-scale C/C++ codebases, making it an optimal choice for modern software security applications.
VI-G Ablation Studies
While our main experiments demonstrate that SecureFalcon achieves strong performance in detecting software vulnerabilities, it is crucial to understand the contribution of individual model components and training configurations. To this end, we performed ablation studies by systematically modifying or removing certain model aspects and evaluating their impact on accuracy and F1-score. The results, summarized in Table XV, provide insights into which factors most influence the model’s performance.
We started with the baseline model configuration described in Table IV, which includes positional embeddings, special token embeddings, twelve attention heads, and a dropout probability of 0.1 for both attention and hidden layers. From this baseline, we disabled positional embeddings, observing a noticeable drop in accuracy and F1 score. This indicates that positional information is essential in capturing code structure and identifying vulnerability patterns.
Removing the special token embeddings—derived from programming language keywords, punctuation, and API calls—led to a notable decrease in accuracy and a slight reduction in the F1 score. These domain-specific tokens provide valuable lexical cues, enabling the model to recognize subtle vulnerability signatures that might otherwise be missed.
Similarly, reducing the number of attention heads from twelve to six diminished the model’s performance, suggesting that the rich representational capacity provided by multiple attention heads is necessary for practical reasoning about complex code patterns. Increasing the dropout rate from 0.1 to 0.2 did not improve results beyond the baseline, only modestly reduced performance. This indicates that while regularization helps prevent overfitting, other factors—such as positional information and specialized tokens—are more critical for overall accuracy and robustness.
In summary, the ablation results highlight the importance of maintaining positional embeddings, preserving domain-specific token embeddings, and ensuring sufficient representational capacity through multiple attention heads.
VI-H Limitations and Future Works
SecureFalcon is currently limited in identifying zero-day vulnerabilities because it relies on known vulnerability patterns and cannot detect previously unseen threats. Additionally, being trained exclusively on C/C++ datasets restricts its ability to analyze other programming languages and prevents it from generating detailed reports that explain the root causes of vulnerabilities within the code. As future work, SecureFalcon will integrate Agentic Retrieval-Augmented Generation (RAG) to address these challenges, enabling AI-driven agents to investigate and exploit potential zero-day vulnerabilities autonomously. To uncover previously unknown vulnerabilities, these agents will leverage diverse data sources, including proprietary databases, real-time threat intelligence, and specialized security tools. Expanding SecureFalcon’s training to encompass multiple programming languages will enhance its versatility and effectiveness across various software environments. Using Agentic RAG, SecureFalcon can dynamically adapt its detection methods, ensuring a more proactive and comprehensive approach to identifying and mitigating emerging security threats.
@@ VI-H Limitations and Future Works
In addition to overcoming zero-day vulnerability detection and expanding language support, SecureFalcon aims to enhance its capabilities by transitioning from classification to generation tasks. This evolution will enable SecureFalcon to provide developers with more detailed and actionable insights. Enhancements include generating comprehensive reports that describe detected vulnerabilities, highlighting affected code snippets, and analyzing potential impacts. Furthermore, SecureFalcon will outline steps to reproduce issues, including environment setup, input data, and execution steps, and offer remediation guidance with fix recommendations and secure coding practices. By fine-tuning SecureFalcon for generation tasks and integrating these features with development tools through Agentic RAG, SecureFalcon will ensure real-time feedback and automated code reviews.
@@ VII Conclusion
Acknowledgments
The authors would like to thank the Technology Innovation Institute for providing the computing resources necessary for training the models used in this research. This work would not have been possible without the institute’s support.